# Primary gravitational waves at high frequencies I: Origin of suppression in the power spectrum

### v1

@@ 6.2 Evolution and PS after the smoother transition
During the radiation dominated era, the rescaled mode function $\mu_{k}(\eta)$ is given by Eq. (2.14) with the functions $m_{k}(\eta)$ and $n_{k}(\eta)$ being given by Eq. (2.16). The coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ are obtained by matching the mode function and its derivative during the smooth linear transition [cf. Eq. (6.3)] to the mode function and its time derivative during the radiation dominated era. The matching is to be carried out at the end of the period of the transition, viz. at $\eta_{\mathrm{tr}}=\eta_{\mathrm{e}}+\Delta\eta$. On carrying out the matching, we find that the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ can be written as
@@
$\tau(\eta_{\mathrm{tr}})$ $\equiv\tau_{\mathrm{tr}}=-\alpha\left[k^{2}-U(\eta_{\mathrm{e}}+\Delta\eta)\right]=-\alpha k^{2}=-y_{\mathrm{e}}^{2}X^{-2/3}\,$ (6.27)
@@
$\beta_{k}^{\mathrm{r}}$ $\simeq\frac{q(q-1)A_{k}}{4y_{\mathrm{e}}^{3}}\mathrm{e}^{i[y_{\mathrm{e}}+\pi(q-1)\pi/2]+iy_{\mathrm{e}}Y}\mathrm{e}^{-i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\left[1-\mathrm{e}^{i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\frac{iX}{q(q-1)}\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\right].$ (6.33b)
We should highlight the fact that the coefficient $\beta^{\mathrm{r}}_{k}$ behaves as $y_{\mathrm{e}}^{-3}$ in the limit $y_{\mathrm{e}}\gg 1$. In other words, $\beta_{k}^{\mathrm{r}}\propto k^{-3}$ in the limit $k\gg k_{\mathrm{e}}$. This should be contrasted with the result we had arrived at earlier in the case of instantaneous transition from inflation to the epoch of radiation domination. Recall that, in the case of instantaneous transition we had obtained that $\beta^{\mathrm{r}}_{k}\propto k^{-2}$ in the limit $k\gg k_{\mathrm{e}}$ [cf. Eq. (3.4b)]. This implies that the smoother transition has modified the scale dependence of the coefficient $\beta_{k}^{\mathrm{r}}$. Such a modification will, in turn change the behavior of the PS at large $k$. It should be clear that the behavior $\beta_{k}^{\mathrm{r}}\propto k^{-2}$ at large $k$ is an artifact of the instantaneous transition from inflation to radiation domination.
@@
$\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}=\frac{2}{3}\frac{y_{\mathrm{e}}^{3}}{X}\left\{\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{3/2}-1\right\}\simeq-\frac{q(q-1)}{X}y_{\mathrm{e}}\,$, (6.34)
@@
Figure 7: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation epoch at the conformal time $\eta=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) in the case of the smoother transition from inflation to radiation domination. As in Figs. 3, 4 and 5, we have set $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ in plotting this figure. Also, we have chosen the width of the smooth transition to be $\Delta\eta=2|\eta_{\mathrm{e}}|$. The effects of the smoother transition from inflation to radiation domination modify the PS over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, we have plotted the actual and regularized PS only over this domain. Clearly, the regularized PS oscillates about zero (over $y_{\mathrm{e}}\gg 1$), as in the case of the instantaneous transition, which we had illustrated in Fig. 4. However, note that, in the case of the smoother transition, the amplitude of the oscillations decreases with $k$ at large wave numbers.
@@
Figure 8: The regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation domination at the conformal time $\eta=2|\eta_{\mathrm{e}}|$, have been plotted (in red and blue) in the cases of instantaneous and smoother transitions from inflation to radiation domination. The nature of the transition from inflation to radiation domination affects the PS only over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, as in the previous figure, we have illustrated the PS only over this domain. Note that, while the regularized PS has a constant amplitude over small scales (such that over $y_{\mathrm{e}}\gg 1$) in the case of the instantaneous transition, clearly, the amplitude of the oscillations decreases as $k^{-1}$ in the case of the smooth linear transition.
@@ 6.3 PS during matter domination, after the smoother transition
In this section, we shall discuss the behavior of PS of PGWs during the epoch of matter domination, after the smooth linear transition from inflation to the epoch of radiation domination. Earlier, we had expressed the actual and the regularized PS of PGWS during matter domination in terms of the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ [see Eqs. (3.2) and (4.2)]. Recall that, these coefficients were, in turn, dependent on the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ during the epoch of radiation domination [see Eqs. (3.21)]. In the case of the smooth linear transition from inflation to radiation domination, we can substitute $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ from Eqs. (6.31) in Eqs. (3.21) to arrive at the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$. The corresponding actual and regularized PS of PGWs during the epoch of matter domination can be arrived at by substituting these expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in Eqs. (3.2) and (4.2). In Fig. 9, we have plotted the actual and regularized PS of PGWs evaluated today. Evidently, the process regularization as well as the smoother transition from inflation to radiation domination affects the PS only over wave numbers such that $k\gtrsim k_{\mathrm{e}}$. Note that the process of regularization truncates the rise in the PS over $k\gtrsim k_{\mathrm{e}}$, as we have already discussed. The regularized PS oscillates about zero over this domain. The smooth linear transition leads to a suppression in the amplitude of the oscillations in the regularized PS (as $k^{-1}$) at such high wave numbers.
Figure 9: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today, i.e. during the late stages of matter domination at the conformal time $\eta_{0}$, have been plotted (in red and blue), assuming a smooth linear transition from de Sitter inflation to the epoch of radiation domination. In plotting the figure, we have assumed that the transition from radiation to matter domination is instantaneous. As in Figs. 3 and 5, the vertical lines indicate the frequencies associated with the wave numbers $(k_{0}/a_{0},k_{\mathrm{eq}}/a_{0},k_{\mathrm{e}}/a_{0})$. We have illustrated the PS only until the frequency of $f\simeq 10^{11}\,\mathrm{Hz}$, since, at higher frequencies, the numerical errors in plotting the regularized PS become significant. Note that the process of regularization does not alter the PS over wave numbers $k\lesssim k_{\mathrm{e}}$. However, as in Fig. 5, it is clear that regularization truncates the $k^{2}$ rise in the PS over $k\gtrsim k_{\mathrm{e}}$. In addition, the smooth linear transition from inflation to radiation domination suppresses the amplitude of the oscillations in the PS over $k\gtrsim k_{\mathrm{e}}$ as $k^{-1}$.
@@
Note that, if we now consider the regularized PS, the first term will not be present. On comparing the last two terms with the second term, we find that, around $k\gtrsim k_{\mathrm{e}}$, the last two terms will dominate. These terms oscillate about zero as in the case of the regularized PS over this domain in the case of the instantaneous transition from inflation to radiation domination [see Eq. (4.20)]. But, because of the smoother transition from inflation to the epoch of radiation domination, the amplitude of the oscillations is suppressed as $k^{-1}$. However, interestingly, we find that the second term eventually becomes dominant in the regularized PS over $k\gtrsim k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$. Therefore, at ultra-high wave numbers, we find that the regularized PS behaves as
@@
The third and last case corresponds to $y_{\mathrm{e}}\ll 1$, $x_{\mathrm{eq}}\ll 1$ and $z_{0}\ll 1$. These limit correspond to wave numbers $k\lesssim k_{0}$, which are yet to enter the Hubble radius. In this domain, the PS during the epoch of matter domination can be written as
@@
Finally, let us point out that, in the limit of instantaneous transition from de Sitter inflation to the epoch of radiation domination, we have $X\rightarrow\infty,\penalty 10000\ Y\rightarrow-1,\penalty 10000\ (\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})\rightarrow 2y_{\mathrm{e}}/X$. In such a limit, we find that $\mathcal{F}\simeq-1/2$ and $\mathcal{G}\simeq 2/(3X^{2})$ and the above expressions (6.49), (6.51), and (6.53) for the PS reduce to the expressions (3.26), (3.28) and (3.29) we had obtained earlier.
@@ 7.1 Summary
In this work, we have examined the effects of regularization and smoothing of the transition from inflation to the epoch of radiation domination on the PS of GWs. We have argued that regularization is essential if we are to avoid the $k^{2}$ rise in the PS on small scales. We have adopted the method of adiabatic subtraction to regularize the PS. We have evolved the PGWs until the time of radiation-matter equality and today, and regularized the PS at these instances. Importantly, we find that regularization does not affect the standard predictions over the wave numbers $k\lesssim k_{\mathrm{e}}$. Moreover, we find that, in the case of instantaneous transition from de Sitter inflation to the epoch of radiation domination, the regularized PS of PGWs does not exhibit a rise over wave numbers such that $k\gtrsim k_{\mathrm{e}}$, i.e. over wave numbers that always remain inside the Hubble radius. Instead, the PS oscillates about zero with a constant amplitude in this domain. We also showed that the regularization of the PS is not sufficient. It is also necessary to smooth the transitions in the cosmological scenario. It is only the combination of regularization and smooth transitions that can lead to well-behaved correlation functions in real space. This is of course natural since realistic transitions are expected to be smooth and, in fact, they are expected to be infinitely continuous. We worked with a linear function to smooth the ‘effective potential’ $U(\eta)$ during the transition from inflation to the epoch of radiation domination. We obtained exact solutions to the mode functions describing the PGWs during the transition and evaluated the PS of GWs at the time of radiation-matter equality as well as today. We showed that, in contrast to the case of instantaneous transition, the smoother transition leads to a suppression in the regularized PS of PGWs over wave numbers $k\gtrsim k_{\mathrm{e}}$. Over such high wave numbers, while the regularized PS continues to oscillate about zero as in the case of the instantaneous transition, it does so with a decreasing amplitude. Further, we encountered an interesting behavior in the regularized PS evaluated today, after a smoother transition from inflation to radiation domination and an instantaneous transition from radiation to matter domination. In this scenario, we found that the suppression in the regularized PS ceases at ultra-high wave numbers such that $k\gg k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$ and the amplitude of the oscillations in the PS becomes constant thereafter. This behavior can be attributed to the assumption of an instantaneous transition from radiation to matter domination. The result clearly suggests that, unless all the intermediate transitions are smoothed, the PS of PGWs evaluated in a given epoch will carry imprints of previous, unrealistic, instantaneous transitions.

### v2

@@ 6.2 Evolution and PS after the smoother transition
During the radiation-dominated era, the rescaled mode function $\mu_{k}(\eta)$ is given by Eq. (2.14), with the functions $m_{k}(\eta)$ and $n_{k}(\eta)$ being given by Eq. (2.16). The coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ are obtained by matching the mode function and its derivative during the smoother linear transition [cf. Eq. (6.3)] with the mode function and its time derivative during the radiation-dominated era. The matching has to be carried out at the end of the period of the transition, viz. at $\eta_{\mathrm{tr}}=\eta_{\mathrm{e}}+\Delta\eta$. On carrying out the matching, we find that the coefficients $(\alpha^{\mathrm{r}}_{k},\beta^{\mathrm{r}}_{k})$ can be written as
@@
$\tau(\eta_{\mathrm{tr}})$ $\equiv\tau_{\mathrm{tr}}=-\alpha\left[k^{2}-U(\eta_{\mathrm{e}}+\Delta\eta)\right]=-\alpha k^{2}=-y_{\mathrm{e}}^{2}X^{-2/3},$ (6.27)
@@
$\beta_{k}^{\mathrm{r}}$ $\simeq\frac{q(q-1)A_{k}}{4y_{\mathrm{e}}^{3}}\mathrm{e}^{i[y_{\mathrm{e}}+(q-1)\pi/2]+iy_{\mathrm{e}}Y}\mathrm{e}^{-i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\left[1-\mathrm{e}^{i(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})}\frac{iX}{q(q-1)}\sin\left(\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}\right)\right].$ (6.33b)
We should highlight the fact that the coefficient $\beta^{\mathrm{r}}_{k}/A_{k}$ behaves as $y_{\mathrm{e}}^{-3}$ in the limit $y_{\mathrm{e}}\gg 1$. In other words, $\beta_{k}^{\mathrm{r}}/A_{k}\propto k^{-3}$ in the limit $k\gg k_{\mathrm{e}}$. This should be contrasted with the result we had arrived at earlier in the case of instantaneous transition from inflation to the epoch of radiation domination. Recall that, in the case of instantaneous transition we had obtained that $\beta^{\mathrm{r}}_{k}/A_{k}\propto k^{-2}$ in the limit $k\gg k_{\mathrm{e}}$ [cf. Eq. (3.4b)]. This implies that the smoother transition has modified the scale dependence of the coefficient $\beta_{k}^{\mathrm{r}}$. Such a modification will, in turn change the behavior of the PS at large $k$. It should be clear that the behavior $\beta_{k}^{\mathrm{r}}/A_{k}\propto k^{-2}$ at large $k$ is an artifact of the instantaneous transition from inflation to radiation domination.
@@
$\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}}=\frac{2}{3}\frac{y_{\mathrm{e}}^{3}}{X}\left\{\left[1-\frac{q(q-1)}{y_{\mathrm{e}}^{2}}\right]^{3/2}-1\right\}\simeq-\frac{q(q-1)}{X}y_{\mathrm{e}}$ (6.34)
@@
Figure 7: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation domination at the conformal time $\eta=2|\eta_{\mathrm{e}}|$ have been plotted (in red and blue) in the case of the smoother transition from inflation to radiation domination. As in Figs. 3, 4 and 5, we have set $g_{\ast,\mathrm{rh}}^{1/4}T_{\mathrm{rh}}=1.8\times 10^{16}\,\mathrm{GeV}$ in plotting this figure. Also, we have chosen the width of the smooth transition to be $\Delta\eta=2|\eta_{\mathrm{e}}|$. The effects of the smoother transition from inflation to radiation domination modify the PS over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, we have plotted the actual and regularized PS only over this domain. Clearly, the regularized PS oscillates about zero (over $y_{\mathrm{e}}\gg 1$), as in the case of the instantaneous transition, which we had illustrated in Fig. 4. However, note that, in the case of the smoother transition, the amplitude of the oscillations decreases with $k$ at large wave numbers.
@@
Figure 8: The regularized PS of PGWs $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated during the early stages of the epoch of radiation domination at the conformal time $\eta=2|\eta_{\mathrm{e}}|$, has been plotted in the cases of instantaneous and smoother transitions from inflation to radiation domination (in red and blue). The nature of the transition from inflation to radiation domination affects the PS only over wave numbers such that $y_{\mathrm{e}}\gtrsim 1$. Hence, as in the previous figure, we have illustrated the PS only over this domain. Note that, while the regularized PS has a constant amplitude over small scales (such that over $y_{\mathrm{e}}\gg 1$) in the case of the instantaneous transition, clearly, the amplitude of the oscillations decreases as $k^{-1}$ in the case of the smoother linear transition.
@@ 6.3 PS during matter domination, after the smoother transition
In this section, we shall discuss the behavior of PS of PGWs during the epoch of matter domination, after the smoother linear transition from inflation to the epoch of radiation domination. Earlier, we had expressed the actual and the regularized PS of PGWS during matter domination in terms of the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ [see Eqs. (3.2) and (4.2)]. Recall that, these coefficients were, in turn, dependent on the Bogoliubov coefficients $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ during the epoch of radiation domination [see Eqs. (3.21)]. In the case of the smoother linear transition from inflation to radiation domination, we can substitute $(\alpha_{k}^{\mathrm{r}},\beta_{k}^{\mathrm{r}})$ from Eqs. (6.31) in Eqs. (3.21) to arrive at the coefficients $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$. The corresponding actual and regularized PS of PGWs during the epoch of matter domination can be arrived at by substituting these expressions for $(\alpha_{k}^{\mathrm{m}},\beta_{k}^{\mathrm{m}})$ in Eqs. (3.2) and (4.2). In Fig. 9, we have plotted the actual and regularized PS of PGWs evaluated today. Evidently, the process regularization as well as the smoother transition from inflation to radiation domination affects the PS only over wave numbers such that $k\gtrsim k_{\mathrm{e}}$. Note that the process of regularization truncates the rise in the PS over $k\gtrsim k_{\mathrm{e}}$, as we have already discussed. The regularized PS oscillates about zero over this domain. The smoother linear transition leads to a suppression in the amplitude of the oscillations in the regularized PS (as $k^{-1}$) at such high wave numbers.
Figure 9: The actual and the regularized PS of PGWs, viz. $\mathcal{P}_{{}_{\mathrm{T}}}(k,\eta)$ and $\mathcal{P}_{{}_{\mathrm{T}}}^{\mathrm{reg}}(k,\eta)$, evaluated today, i.e. during the late stages of matter domination at the conformal time $\eta_{0}$, have been plotted (in red and blue), assuming a smoother linear transition from de Sitter inflation to the epoch of radiation domination. In plotting the figure, we have assumed that the transition from radiation to matter domination is instantaneous. As in Figs. 3 and 5, the vertical lines indicate the frequencies associated with the wave numbers $(k_{0}/a_{0},k_{\mathrm{eq}}/a_{0},k_{\mathrm{e}}/a_{0})$. We have illustrated the PS only until the frequency of $f\simeq 10^{11}\,\mathrm{Hz}$, since, at higher frequencies, the numerical errors in plotting the regularized PS become significant. Note that the process of regularization does not alter the PS over wave numbers $k\lesssim k_{\mathrm{e}}$. However, as in Fig. 5, it is clear that regularization truncates the $k^{2}$ rise in the PS over $k\gtrsim k_{\mathrm{e}}$. In addition, the smoother linear transition from inflation to radiation domination suppresses the amplitude of the oscillations in the PS over $k\gtrsim k_{\mathrm{e}}$ as $k^{-1}$.
@@
Note that, if we now consider the regularized PS, the first term will not be present. On comparing the last two terms with the second term, we find that, around $k\gtrsim k_{\mathrm{e}}$, the last two terms will dominate. These terms oscillate about zero as in the regularized PS over this domain in the case of the instantaneous transition from inflation to radiation domination [see Eq. (4.20)]. But, because of the smoother transition from inflation to the epoch of radiation domination, the amplitude of the oscillations is suppressed as $k^{-1}$. However, interestingly, we find that the second term eventually becomes dominant in the regularized PS over $k\gtrsim k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$. Therefore, at ultra-high wave numbers, we find that the regularized PS behaves as
@@
The third and last case corresponds to $y_{\mathrm{e}}\ll 1$, $x_{\mathrm{eq}}\ll 1$ and $z_{0}\ll 1$. These limit correspond to wave numbers $k\lesssim k_{0}$, which are yet to re-enter the Hubble radius. In this domain, the PS during the epoch of matter domination can be written as
@@
Finally, let us point out that, in the limit of instantaneous transition from de Sitter inflation to the epoch of radiation domination, we have $X\rightarrow\infty,\penalty 10000\ Y\rightarrow-1,\penalty 10000\ (\zeta_{\mathrm{e}}-\zeta_{\mathrm{tr}})\rightarrow 2y_{\mathrm{e}}/X$. In such a limit, we find that $\mathcal{F}\simeq-1/2$ and $\mathcal{G}\simeq 2/(3X^{2})$ and the above expressions (6.49), (6.51) and (6.53) for the PS reduce to the expressions (3.26), (3.28) and (3.29) we had obtained earlier.
@@ 7.1 Summary
In this work, we have examined the effects of regularization and smoothing of the transition from inflation to the epoch of radiation domination on the PS of GWs. We have argued that regularization is essential if we are to avoid the $k^{2}$ rise in the PS on small scales. We have adopted the method of adiabatic subtraction to regularize the PS. We have evolved the PGWs until the time of radiation-matter equality and today, and regularized the PS at these instances. Importantly, we find that regularization does not affect the standard predictions over the wave numbers $k\lesssim k_{\mathrm{e}}$. Moreover, we find that, in the case of instantaneous transitions from de Sitter inflation to the epochs of radiation and matter domination, the regularized PS of PGWs does not exhibit a rise over wave numbers such that $k\gtrsim k_{\mathrm{e}}$, i.e. over wave numbers that always remain inside the Hubble radius. Instead, the PS oscillates about zero with a constant amplitude in this domain.
We also showed that, when transitions are involved in the cosmological scenarios of interest, it becomes important to take into account the effects due to the smoothness of the transitions. We should clarify that the process of regularization and smoothing of the transitions play distinct roles in the determining structure of the primordial correlation functions. It seems reasonable to expect that, across transitions, the combination of regularization and smooth transitions can ensure that generic correlation functions in real space are well-behaved. This is of course natural since realistic transitions are expected to be smooth and, in fact, they are expected to be infinitely continuous. We worked with a linear function to smooth the ‘effective potential’ $U(\eta)$ during the transition from inflation to the epoch of radiation domination. We obtained exact solutions to the mode functions describing the PGWs during the transition and evaluated the PS of GWs at the time of radiation-matter equality as well as today. We showed that, in contrast to the case of instantaneous transition, the smoother transition leads to a suppression in the regularized PS of PGWs over wave numbers $k\gtrsim k_{\mathrm{e}}$. Over such high wave numbers, while the regularized PS continues to oscillate about zero as in the case of the instantaneous transition, it does so with a decreasing amplitude. Further, we encountered an interesting behavior in the regularized PS evaluated today, after a smoother transition from inflation to radiation domination and an instantaneous transition from radiation to matter domination. In this scenario, we found that the suppression in the regularized PS ceases at ultra-high wave numbers such that $k\gtrsim k_{\mathrm{e}}^{3}/k_{\mathrm{eq}}^{2}$ and the amplitude of the oscillations in the PS becomes constant thereafter. This behavior can be attributed to the assumption of an instantaneous transition from radiation to matter domination. The result clearly suggests that, unless all the intermediate transitions are completely smoothed, the PS of PGWs evaluated in a given epoch will carry imprints of earlier, unrealistic, instantaneous transitions.